# What You See Is Not Always What You Get: An Empirical Study of Code Comprehension by Large Language Models

### v1

@@ 5. Results
Perturbation budget Confidence% Correctness%
Clean 95.47 100
20% 54.81 72.67
40% 38.42 54.67
60% 26.91 47.33
80% 11.28 27.67
100% 3.62 16.33
Table 2. Reordering
Perturbation budget Confidence% Correctness%
Clean 96.11 100
20% 51.80 65.67
40% 42.85 58.00
60% 35.15 50.67
80% 33.36 51.00
100% 29.64 47.67
Table 3. Invisible characters
Perturbation budget Confidence% Correctness%
Clean 95.45 100
20% 41.84 60.33
40% 20.65 42.67
60% 10.11 27.33
80% 7.42 24.33
100% 1.74 15.33
Table 4. Deletions
Perturbation budget Confidence% Correctness%
Clean 95.74 100
20% 77.85 86.67
40% 58.47 69.33
60% 49.93 63.00
80% 51.33 68.00
100% 54.23 67.67
Table 5. Homoglyphs
From the Tables 2 - 5, we can immediately find many similarities across the 4 categories. As expected, the confidence score for the clean code snippet is high with none below 95%, and the correctness is 100% across the board. The highest confidence score is held by invisible characters at 96.11%. Each category has their largest drop in confidence between clean and 20% perturbation budget, with differences getting smaller as the perturbation budget rises. The largest difference between budget levels is held by the deletions category between clean and 20%, at -52.76%. Reorderings and deletions have significantly lower floors for both confidence and correctness scores compared to invisible characters and homoglyphs, both approaching 0% for confidence, and both sub-20% for correctness. Deletions holds the lowest score across all perturbation budget levels for both confidence at 1.74% as well as correctness at 15.3%. The highest scores at 100% perturbation budget are held by the homoglyph category, with confidence at 54.23% and correctness at 67.67%. The homoglyph category is also the only perturbation category to not show decreases in scores at every level, with an increase in both confidence and correctness scores recorded between 60%-100%.
@@
Perturbation budget Confidence% Correctness%
Clean 93.51 100
20% 28.55 46.33
40% 18.80 36.33
60% 15.94 34.33
80% 11.83 26.33
100% 8.44 22.00
Table 6. Reordering
Perturbation budget Confidence% Correctness%
Clean 93.84 100
20% 48.61 66.33
40% 44.16 62.67
60% 35.71 54.67
80% 31.03 48.33
100% 37.38 56.67
Table 7. Invisible characters
Perturbation budget Confidence% Correctness%
Clean 93.69 100
20% 19.02 34.67
40% 8.05 19.00
60% 3.93 17.00
80% 4.80 13.67
100% 4.60 11.33
Table 8. Deletions
Perturbation budget Confidence% Correctness%
Clean 93.46 100
20% 75.16 84.00
40% 58.91 72.67
60% 49.71 64.33
80% 43.94 60.67
100% 49.80 64.33
Table 9. Homoglyphs
@@ 5.3. GPT-4o-2024-05-13
Perturbation budget Confidence% Correctness%
Clean 80.92 100
20% -18.35 0.67
40% -19.08 0
60% -18.59 0.67
80% -18.82 0.33
100% -17.24 2.33
Table 10. Reordering
Perturbation budget Confidence% Correctness%
Clean 81.46 100
20% -17.05 1.33
40% -16.34 2.00
60% -17.43 1.00
80% -17.80 0.67
100% -18.54 0
Table 11. Invisible characters
Perturbation budget Confidence% Correctness%
Clean 80.94 100
20% -18.29 0.67
40% -18.31 0.67
60% -18.27 1.00
80% -17.92 1.00
100% -17.52 1.33
Table 12. Deletions
Perturbation budget Confidence% Correctness%
Clean 81.28 100
20% -17.54 1.00
40% -17.58 1.00
60% -18.00 0.67
80% -18.35 0.33
100% -18.36 0.33
Table 13. Homoglyphs
Tables 10 - 13 display a clear and significant difference to the two GPT-3.5 models. However, there is one characteristic held in common across the results of all three models. Across all 4 categories, the clean code is still 100% correct and still holds the highest confidence score, with all of them holding a score of 81±1%. However, from there, things change drastically. The GPT-4 results are the only results to show negative confidence scores, which is due to the disproportionate number of incorrect answers, with the lowest score of -19.08% held by reorders at 20% perturbation budget. As with confidence scores, correctness is also extremely low across the board, with the next highest score following 100% being 2.33%, also held by reorders at 100% perturbation budget. The lowest correctness score is tied between reorders at 20% perturbation budget and invisible characters at 100% perturbation budget, and is also the lowest possible correctness, 0%.
5.4. Summary
The similarities between the two GPT-3.5 models and their difference compared to the GPT-4o model reveal nuanced insights into the impact of perturbations on their respective performance outcomes. The GPT-3.5 models exhibit uniform decline in both confidence and correctness, with some categories exhibiting a natural limit when approaching 100% perturbation. In contrast to that, the GPT-4o model shows a massive decrease in both confidence and correctness at any perturbation budget level, with correctness approaching 0 across all categories. This seems to indicate that the newer model has some kind of special character detection and vetting that the older GPT-3.5 generation models do not have.
Critically, all models responded with 100% correctness for the clean code, and all models demonstrated a significant decline in both scores with any amount of perturbation. Whilst evaluating code comprehension was not an explicit experiment objective, it is helpful for the overall research goals to see data that validates the baseline code comprehension skills of these large language models.
@@ 5.5. Answer to research questions
RQ1: To what extent do imperceptible perturbations within code correlate with LLM performance outcomes in software testing scenarios?
@@
Taking another look at the raw data explains why this is the case.
@@
For reference, Tables. 16 and 17 show the proportion of total correct responses to incorrect responses returned by the two GPT-3.5 models. It is immediately obvious that for these models, there is a statistically significant number of correct responses for each category.
@@
RQ2: Does a correlation exist between imperceptible perturbations and LLM ‘confidence’, and what implications does this have for software testing?
@@
RQ3: How significantly do different imperceptible perturbation methods impact the responses of an LLM model?
@@
RQ4: How do advancements in recent LLM iterations impact their robustness against imperceptible adversarial perturbations in software testing scenarios?
A trend throughout the results suggests significant internal differences between the GPT-3.5 generation models and the GPT-4 generation model. We can deduce that some deterministic guardrails have been put in place for the GPT-4 generation model which the GPT-3.5 generation models do not have. Whilst the GPT-3.5 models lose confidence and correctness proportionally to the amount of perturbation present, GPT-4o returns an incorrect response across the board when presented with any category of perturbation, at any budget level.
@@ 7. Conclusion
The findings in this paper indicate that where some LLMs were susceptible to imperceptible attacks, some newer version of LLMs have enacted measures to guard against these false positives, such ChatGPT4o. However, even between the older 3.5 generation models, the newer model showed marked improvement in comparison to the older model. Throughout the study, we observed that perturbation budget had a strong negative correlation to performance outcomes for code analysis and comprehension task. There is a strong negative correlation between the presence of perturbation and performance outcomes for newer version of LLMs, but no valid correlational relationship between perturbation budget and performance outcomes. We claim that the current devised solutions of imperceptible character attacks have presented significant impacts for the coding tasks. We anticipate future research will focus on bridging such issue to enhance the trust of LLMs, in particular for code analysis and comprehension tasks.

### v2

@@ 5.1. Overview
From the Table 2, we can immediately find many similarities across the 4 categories. As expected, the confidence score for the clean code snippet is high with none below 95%, and the correctness is 100% across the board. The highest confidence score is held by invisible characters at 96.11%. Each category has their largest drop in confidence between clean and 20% perturbation budget, with differences getting smaller as the perturbation budget rises. The largest difference between budget levels is held by the deletions category between clean and 20%, at -52.76%. Reorderings and deletions have significantly lower floors for both confidence and correctness scores compared to invisible characters and homoglyphs, both approaching 0% for confidence, and both sub-20% for correctness. Deletions holds the lowest score across all perturbation budget levels for both confidence at 1.74% as well as correctness at 15.3%. The highest scores at 100% perturbation budget are held by the homoglyph category, with confidence at 54.23% and correctness at 67.67%. The homoglyph category is also the only perturbation category to not show decreases in scores at every level, with an increase in both confidence and correctness scores recorded between 60%-100%.
Perturbation budget Reordering Invisible characters Deletions Homoglyphs
Conf.% Corr.% Conf.% Corr.% Conf.% Corr.% Conf.% Corr.%
Clean 95.47 100 96.11 100 95.45 100 95.74 100
20% 54.81 72.67 51.80 65.67 41.84 60.33 77.85 86.67
40% 38.42 54.67 42.85 58.00 20.65 42.67 58.47 69.33
60% 26.91 47.33 35.15 50.67 10.11 27.33 49.93 63.00
80% 11.28 27.67 33.36 51.00 7.42 24.33 51.33 68.00
100% 3.62 16.33 29.64 47.67 1.74 15.33 54.23 67.67
Table 2. Performance of the model GPT-3.5-turbo-0613 under four imperceptible coding character attack methods.
@@
Perturbation budget Reordering Invisible characters Deletions Homoglyphs
Conf.% Corr.% Conf.% Corr.% Conf.% Corr.% Conf.% Corr.%
Clean 93.51 100 93.84 100 93.69 100 93.46 100
20% 28.55 46.33 48.61 66.33 19.02 34.67 75.16 84.00
40% 18.80 36.33 44.16 62.67 8.05 19.00 58.91 72.67
60% 15.94 34.33 35.71 54.67 3.93 17.00 49.71 64.33
80% 11.83 26.33 31.03 48.33 4.80 13.67 43.94 60.67
100% 8.44 22.00 37.38 56.67 4.60 11.33 49.80 64.33
Table 3. Performance of the model GPT-3.5-turbo-0125 under four imperceptible coding character attack methods.
@@ 5.4. GPT-4o-2024-05-13
Table 4 displays a clear and significant difference to the two GPT-3.5 models. However, there is one characteristic held in common across the results of all three models. Across all 4 categories, the clean code is still 100% correct and still holds the highest confidence score, with all of them holding a score of 81±1%. However, from there, things change drastically. The GPT-4o results are the only results to show negative confidence scores, which is due to the disproportionate number of incorrect answers, with the lowest score of -19.08% held by reorders at 20% perturbation budget. As with confidence scores, correctness is also extremely low across the board, with the next highest score following 100% being 2.33%, also held by reorders at 100% perturbation budget. The lowest correctness score is tied between reorders at 20% perturbation budget and invisible characters at 100% perturbation budget, and is also the lowest possible correctness, 0%.
Perturbation budget Reordering Invisible characters Deletions Homoglyphs
Conf.% Corr.% Conf.% Corr.% Conf.% Corr.% Conf.% Corr.%
Clean 80.92 100 81.46 100 80.94 100 81.28 100
20% -18.35 0.67 -17.05 1.33 -18.29 0.67 -17.54 1.00
40% -19.08 0 -16.34 2.00 -18.31 0.67 -17.58 1.00
60% -18.59 0.67 -17.43 1.00 -18.27 1.00 -18.00 0.67
80% -18.82 0.33 -17.80 0.67 -17.92 1.00 -18.35 0.33
100% -17.24 2.33 -18.54 0 -17.52 1.33 -18.36 0.33
Table 4. Performance of the model GPT-4o-2024-05-13 under four imperceptible coding character attack methods.
@@ 5.5. Answer to research questions
RQ1: To what extent do imperceptible perturbations within code correlate with LLMs’ performance outcomes in software testing scenarios?
@@
For reference, Tables. 7 and 8 show the proportion of total correct responses to incorrect responses returned by the two GPT-3.5 models. It is immediately obvious that for these models, there is a statistically significant number of correct responses for each category.
@@
RQ2: Does a correlation exist between imperceptible perturbations and LLMs’ ‘confidence’, and what implications does this have for software testing?
@@
RQ3: How significantly do different imperceptible perturbation methods impact the responses of LLMs?
@@
RQ4: How do advancements in recent LLMs development improve their robustness against imperceptible adversarial perturbations in software testing scenarios?
A trend throughout the results suggests significant internal differences between the GPT-3.5 generation models and the GPT-4o generation model. We can deduce that some deterministic guardrails have been put in place for the GPT-4o generation model which the GPT-3.5 generation models do not have. Whilst the GPT-3.5 models lose confidence and correctness proportionally to the amount of perturbation present, GPT-4o returns an incorrect response across the board when presented with any category of perturbation, at any budget level.
@@ 7. Conclusion
The findings in this paper identify the vulnerabilities in recent LLMs for imperceptible coding character attack. With the black-box imperceptible coding character attack framework, we have introduced four distinct attack strategies that effectively manipulate the LLMs’ behaviour in coding related tasks. We have conducted large-scale experiments to demonstrate the susceptibility of LLMs against code snippet perturbations. Furthermore, we have observed that perturbation budget had a strong negative correlation to performance outcomes for code analysis and comprehension task. Additionally, concerning the latest LLMs solution, there is a strong negative correlation between the presence of perturbation and performance outcomes, but no valid correlational relationship between perturbation budget and performance outcomes. We claim that the current devised solutions of imperceptible coding character attacks have presented significant impacts for the coding tasks. We anticipate future research will focus on bridging such issue to enhance the trust of LLMs, in particular for code analysis and comprehension tasks.